# Gravitational odd-parity perturbation of a Horndeski hairy black hole: quasinormal mode and parameter constraint

### v2

@@ Abstract
In the binary black hole coalescence, the gravitational wave emitted at the ringdown stage can be well described within the black hole perturbation theory, where the quasinormal modes (QNMs) become the important ingredient in modeling the pattern wave form. In general ralativity (GR), the QNMs can be obtained from solving the Regge-Wheeler equation in static black hole, while in Horndeski gravity, the metric perturbation equation can be simplified into a modified Regge-Wheeler equation from the perturbed action. In this paper, we calculate the QNMs frequencies of the gravitational odd-parity perturbation of a specific hairy black hole in Horndeski gravity with the use of the matrix method and pseudo spectral method. Our results indicate that such a Horndeski hairy black hole is stable under the odd perturbation, which is also verified by the time evolution of the perturbation. In particular, we find that for a certain range of the Horndeski hair, the $\ell>2$ modes become the long-lived mode instead of $\ell=2$ mode in GR. Then, we use the ringdown QNMs to preliminarily investigate the signal-to-noise-ratio (SNR) rescaled measurement error of the Horndeski hair. We obtained significant effects of the angular momentum and overtone on the error bound of the hairy parameter. We hope that our findings could inspire more theoretical and phenomenal work on the test of no-hair theorem of black hole from gravitational wave physics.
@@ I Introduction
Testifying the no hair theorem of black hole is important to test the nature of gravity and more fundamental physics, which inevitably requires us to explore the strong gravity regime. The potential ways of seeking the black hole hair are to investigate the imprints of hair on the gravitaional wave (GW) and on the shadow from the black hole. Recent GW observations from the merger of black holes or other compact objects Abbott et al. (2016, 2019, 2020) and black hole shadows from supermassive black holes Akiyama et al. (2019, 2022) open new era of exploring physics about strong gravity regime, which trigger extensive investigation about the prints of hair on the GW and shadows. The studies on testing the no hair theorem with Event Horizon Telescope (EHT) observations can refer to, for examples, the literatures Broderick et al. (2014); Psaltis et al. (2016); Khodadi et al. (2020, 2021); Tang and Xu (2022), but in this scenario, a universal challenge is the high precision in detecting the print of hair on black hole shadow. This is because the hair should directly affect the orbits very near the black hole horizon, but the photon ring is still far away from the horizon. GWs generated from extreme mass ratio inspiral (EMRI) and ringdown stage in binary black hole coalescence are potential tools to seek the black hole hair Thrane et al. (2017); Isi et al. (2019). In the stage of EMRI the hair will influence the inspiring orbits and energy radiation Rodriguez et al. (2012); Kuntz et al. (2020); Maselli et al. (2020); Collodel et al. (2022); Zi et al. (2021); Guo et al. (2022); Liang et al. (2023); Zhang et al. (2023); Zi and Li (2023); Guo et al. (2024), while in ringdown stage, the hair will characterize the ringdown GW wave Kamaretsos et al. (2012); Gossan et al. (2012); Meidam et al. (2014); Shi et al. (2019); Bhagwat et al. (2020). Therefore, it is important to concern both cases to completely trace the imprint of black hole hair on the GWs.
@@
Introducing scalar field into the action of a gravitational sector is a natural way to construct black hole with scalar hair. Among the numerous gravity theories, the Horndeski gravity containing higher derivatives of a scalar field $\phi$ and metric tensor $g_{\mu\nu}$ have retained lots of attentions, since it possesses at most second-order differential field equations so as to be free of Ostrogradski instabilities Horndeski (1974). Various observational constraints or bounds on Horndeski theories have been explored in Bellini et al. (2016); Bhattacharya and Chakraborty (2017); Kreisch and Komatsu (2018); Hou and Gong (2018); Spurio Mancini et al. (2019); Allahyari et al. (2020); Ghosh et al. (2023) and references therein. Horndeski gravity is also active in the cosmological and astrophysical communities since it has significant consequences in describing the accelerated expansion and other interesting features Kobayashi (2019) . Moreover, since Horndeski theory contains a scalar field, and similar to GR it has diffeomorphism invariance and second-order field equations, so it also provides a natural platform to testify the no-hair theorem of black hole. In fact, hairy black holes in Horndeski gravity have been widely studied, for instances, the radially dependent scalar field Rinaldi (2012); Cisterna and Erices (2014); Feng et al. (2015); Sotiriou and Zhou (2014); Miao and Xu (2016); Kuang and Papantonopoulos (2016); Babichev et al. (2016); Benkel et al. (2017); Filios et al. (2019); Cisterna et al. (2018); Giusti et al. (2022); Cisterna et al. (2015); Anabalon et al. (2014); Fontana et al. (2019) and the time-dependent scalar field Babichev and Charmousis (2014); Babichev et al. (2018); Ben Achour and Liu (2019); Takahashi et al. (2019); Minamitsuji and Edholm (2019); Arkani-Hamed et al. (2004). The hairy solution with a scalar hair in linear time dependence was initially ruled out due to instability Khoury et al. (2020), however, another novel solution with a time-dependent scalar hair was recently constructed Bakopoulos et al. (2024) and found to be stable Sirera and Noller (2024). The stability conditions for a similar time-dependent background solution beyond Horndeski theories was also addressed in Mironov et al. (2024). Moreover, it was demonstrated in Hui and Nicolis (2013) , that the no-hair theorem cannot be hold when a Galileon field is coupled to gravity, but the static spherical solution only admits trivial Galileon profile.
@@
An interesting new development in the examination of no-hair theorem in Horndeski theories is that the shift-symmetric Horndeski theory and beyond with the action Babichev et al. (2017)
@@
allow static and asymptotically flat black holes with a nontrivial static scalar field. In the above action, $g$ is the determinant of the metric and $R$ is the Ricci scalar. $\chi=-\partial^{\mu}\phi\partial_{\mu}\phi/2$ is the canonical kinetic term, $Q_{i}~{}(i=2,3,4,5)$ are arbitrary functions of the scalar field $\phi$ and the kinetic term $\chi$ and $Q_{i,\chi}\equiv\partial Q_{i}/\partial\chi$. In particular, a static hairy black hole in a specific quartic Horndeski theory, which includes an additional $log$-term deforming from the standard Schwarzschild black hole, has been constructed in Bergliaffa et al. (2021). This Horndeski hairy black hole soon attracts plenty of attentions, and various theoretical and observational phenomena have been disclosed, for instances, thermodynamic and weak gravitational lensing Walia et al. (2022); Atamurotov et al. (2022), the strong gravitational lensing Kumar et al. (2022), shadow constraint from EHT observation Afrin and Ghosh (2022); Vagnozzi et al. (2023), superradiant properties in its rotating counterpart Jha et al. (2022); Lei et al. (2024), photon rings in the black hole image Wang et al. (2023); Hu et al. (2023) and precessing and periodic orbits Lin and Deng (2023). These studies disclosed the novel properties brought by the Horndeski hair, differentiating from the Schwarzschild black hole in GR, which are helpful to understand the gravitational structure of the specific theory.
@@
Another significant question concerning this Horndeski hairy black hole is its stability. It was addressed in Walia et al. (2022) that the Horndeski hairy black hole is thermodynamically preferred. Later, we also disclosed its dynamical stability under various external massless perturbed fields including the scalar field, electromagnetic field and Dirac field in Yang et al. (2024). However, up to now, its dynamical stability has not been completely investigated. To promote this topic, we have to perturb the metric itself and study the linear mode as a first important step, in which the modes of a spherically symmetric black holes usually can be classified into the even-parity (polar) modes and odd-parity (axial) modes.
@@
On the other hand, considering that the gravitational waves radiated from black hole coalescence contain rich information about the nature of spacetime, it is significant to study the gravitational wave signatures of such hairy black hole. In particular, the ringdown wave, as the associated gravitational emission at the late time of binary black hole coalescence, is expected to be modeled by the superposition of exponentially damped sinusoidal signals. In this stage, the QNMs could be the key observable in constructing the patterns of such signal and therefore could be used to examine the no-hair theorem due to its highly dependency on the black hole parameter. Thus, the first step in this direction is to study their QNMs, which describes the ringdown stage after the binary black holes merger Berti et al. (2015). The QNMs are the eigenmodes of a dissipative system, which inherently reflect the influence of all the system’s parameters. Consequently, if a black hole acquires an additional Horndeski hair, its QNMs will naturally encode the characteristics of this scalar hair. Due to the dissipative nature of the system, the frequencies of QNMs are usually have two parts: the real part and the imaginary part, the former of which gives the frequency of vibration while the latter describes the damping timescale. The readers can refer to the reviews Nollert (1999); Berti et al. (2009); Konoplya and Zhidenko (2011) for more details.
@@
Thus, motivated from the above aspects, in this paper we will consider the dynamic of the gravitational odd-parity perturbations of a specific Horndeski hairy black hole. We expect that it has a simpler spectrum than the even-parity modes because they do not contain the perturbations of the scalar field in contrast to the latter. We shall study the time domain evolution and QNMs frequencies for the odd-parity perturbations, then further answer the question whether the hairy black hole is stable under such perturbations. Moreover, in order to check the imprint of Horndeski hair on the GW in ringdown stage and give some insight into the test of no hair theorem, we will use the yielding QNMs data to construct the ringdown GW wave form and further calculate the Fisher information matrix (FIM) Berti et al. (2006a) to obtain the measurement error of the detection of the Horndeski hair.
@@
This paper is organized as follows. In section II, we derive the modified Regge-Wheeler equation in a specific hairy black hole. In section III, we numerically solve the master equation, and extract the QNMs frequencies of the gravitational odd perturbation. We analyze the effects of scalar hair $Q$, the angular number $\ell$, and the overtones $n$ on the QNM frequencies. In section IV, we first briefly review the waveform modeling in the ringdown stage, and then investigate the detectability of the specific Horndeski hair by calculating the error bound in cases with single-mode and mixed-mode estimations. Section V concludes with our key findings and outlooks for future work. Throughout the paper, we will set $c=\hbar=G=1$ unless we restate it.
@@ II.1 Review on the action and Horndeski hairy black hole
Horndeski gravity is the most general scalar-tensor field theory in four dimensions which possesses second-order field equations Horndeski (1974). In these subsection, we shall focus on a special subclass of Horndeski theory and review the static hairy black hole constructed in Bergliaffa et al. (2021). The action is
@@
and then solving the field equation gives
@@
where $M$ and $Q$ are integral constants. It is noticed from (10) that $\alpha_{42}$ and $\alpha_{22}$ should have different signs to guarantee that the scalar field is real.
Thus, the subclass Horndeski theory with the action (2) with some specific settings admits a Horndeski hairy black hole with the metric Bergliaffa et al. (2021)
@@
in which the hairy term appears as a $log$ term deforming the standard Schwarzschild black hole. Similarly, the integral constants $M$ and $Q$ are related with the mass and hairy charge of the black hole. It is obvious that the metric is asymptotically flat since $f(r)\mid_{r\to\infty}=1$. Let us briefly analyze the structure of the hairy metric (11). For arbitrary $Q$, $r=0$ is an intrinsic singularity as the curvature scalar is singular, and $f(r)=0$ always admits the solution $r_{+}=2M$. In addition, (i) when $Q\to 0$, the metric reduces to Schwarzschild case. (ii) When $Q>0$, $r_{+}=2M$ is the unique root to $f(r)=0$, so it has a unique horizon, i.e., the event horizon for the hairy black hole which resemble Schwarzschild black hole. (iii) While for $-2M<Q<0$, $f(r)=0$ gives two distinguishable positive real roots: the larger one is $r_{+}=2M$ denoting the position of event horizon, and the smaller one $r_{-}$ denotes the Cauchy horizon which is smaller than the event horizon. The Cauchy horizon $r_{-}$ increases as $Q$ decreases, and finally coincides with $r_{+}$ when $Q=-2M$, indicating the extreme case.
@@ II.2 Odd-parity perturbation and modified Regge-Wheeler equation
In this subsection, we will consider the gravitational perturbation to the Horndeski hairy black hole (11). We will focus on the odd-parity perturbation in linear order, where we can ignore the coupling between the even-parity metric perturbation and the perturbation of energy momentum tensor from the scalar field, because both of them decouple from the odd-parity perturbation. The gravitational perturbation of general spherically symmetric black holes in Horndeski gravity was well addressed in Kobayashi et al. (2012, 2014), based on which we shall derive the modified Regge-Wheeler equation for the odd-parity perturbation of the current hairy black hole.
Under the Regge-Wheeler gauge Regge and Wheeler (1957), the odd perturbation to the background metric is
$h_{\mu\nu,\ell m}^{\text{odd}}=$ $\begin{pmatrix}0&0&0&h_{0}(r,t)\\ 0&0&0&h_{1}(r,t)\\ 0&0&0&0\\ h_{0}(r,t)&h_{1}(r,t)&0&0\end{pmatrix}\sin{\theta}\partial_{\theta}Y_{\ell m}$ (12)
where the spherical harmonics can be replaced by Legendre polynomials by setting the azmuthal number $m=0$ without loss of generality, i.e. $Y_{\ell m}\mid_{m=0}=\sqrt{\frac{2\ell+1}{4\pi}}P_{\ell}(\cos\theta)$, because the background hairy metric is spherical symmetric. By substituting the Horndeski hairy metric (11) and the perturbation (12) into the action (2), we can obtain a second order action as

### v3

@@ Abstract
During the binary black hole coalescence, gravitational waves emitted at the ringdown stage can be well described by black hole perturbation theory, where the quasinormal modes (QNMs) become the important ingredient in modeling the pattern waveform. In general relativity (GR), the QNMs can be obtained from solving the Regge-Wheeler (RW) equation of a non-rotating black hole. While in Horndeski gravity, the isospectrality between the odd and even parity perturbations is broken due to the scalar field, but the odd perturbation equation can be simplified into a modified RW equation from the perturbed action. In this paper, we propose a new auxiliary field and tortoise coordinate to refine the modified RW equation in Horndeski gravity, and calculate the QNMs frequencies of the odd perturbation of a specific hairy black hole. It is found that this proposal not only cures the superluminal propagation addressed in the previous literature, but also hold the original QNM spectrum of the odd perturbation. Moreover, our results indicate that such a Horndeski hairy black hole is stable under the odd perturbation, which is also verified by the time evolution of the perturbation. In particular, in contrary to GR, the modes with $\ell=2$ can decay faster than modes with $\ell>2$ for a certain range of the Horndeski hair, and the link between the null geodesics and QNM for the odd perturbation in the current theory is violated. Then, we use the ringdown QNMs to preliminarily investigate the signal-to-noise-ratio (SNR) rescaled measurement error of the Horndeski hair. We obtain significant effects of the angular momentum and overtone on the error bound of the hair parameter. We hope that our findings could inspire more theoretical and phenomenological work on the test of the no-hair theorem of black hole using gravitational wave physics.
@@ I Introduction
Testifying the no-hair theorem of black hole PhysRev.164.1776 ; PhysRevD.5.1239 ; BEKENSTEIN1974535 ; Hawking1972 and its extensions to other theories of gravity is important to test the nature of gravity and more fundamental physics, which inevitably requires us to explore the strong gravity regime. Potential ways of seeking the black hole hair are to investigate the imprints of hair on the gravitaional wave (GW) and on the shadow from the black hole. Recent GW observations from the merger of black holes or other compact objects LIGOScientific:2016aoc ; LIGOScientific:2018mvr ; LIGOScientific:2020aai and black hole shadows from supermassive black holes EventHorizonTelescope:2019dse ; EventHorizonTelescope:2022xnr open a new era of exploring physics about the strong gravity regime, which triggers extensive investigation about the prints of hair on the GW and shadows. The studies on testing the no-hair theorem with Event Horizon Telescope (EHT) observations can refer to, for examples, the literature Broderick:2013rlq ; Psaltis:2015uza ; Khodadi:2020jij ; Khodadi:2021gbc ; Tang:2022uwi , but in this scenario, a universal challenge is achieving high precision in detecting the print of hair on black hole shadow. This is because the hair should directly affect the orbits very near the black hole horizon, but the photon ring still lies far away from the horizon. GWs generated during the extreme mass ratio inspiral (EMRI) and ringdown stage in binary black hole coalescence are potential tools to seek the black hole hair Thrane:2017lqn ; Isi:2019aib . During the stage of EMRI the hair will influence the inspiraling orbits and energy radiation Rodriguez:2011aa ; Kuntz:2020yow ; Maselli:2020zgv ; Collodel:2021jwi ; Zi:2021pdp ; Guo:2022euk ; Liang:2022gdk ; Zhang:2022rfr ; Zi:2023omh ; Guo:2023mhq , while at the ringdown stage, the hair will characterize the ringdown GW wave Kamaretsos:2011um ; Gossan:2011ha ; Meidam:2014jpa ; Shi:2019hqa ; Bhagwat:2019dtm . Therefore, it is important to concern both cases to completely trace the imprint of black hole hair on the GWs.
@@
Introducing a scalar field into the action of a gravitational sector is a natural way to construct black hole with scalar hair. Among the numerous gravity theories, the Horndeski gravity, which contains higher derivatives of a scalar field $\phi$ and metric tensor $g_{\mu\nu}$, has retained lots of attention, since it possesses at most second-order differential field equations so as to be free of Ostrogradski instabilities Horndeski:1974wa . Various observational constraints and bounds on Horndeski theories have been explored in Bellini:2015xja ; Bhattacharya:2016naa ; Kreisch:2017uet ; Hou:2017cjy ; SpurioMancini:2019rxy ; Allahyari:2020jkn ; Ghosh:2023kge and references therein. Horndeski gravity is also actively studied in the cosmological and astrophysical communities since it has significant consequences in describing the accelerated expansion and other interesting features Kobayashi:2019hrl . Moreover, since Horndeski theory contains a scalar field, and similar to GR it has diffeomorphism invariance and second-order field equations, so it also provides a natural platform to verify the no-hair theorem of black hole. In fact, hairy black holes in Horndeski gravity have been widely studied, for instance, the radially dependent scalar field Rinaldi:2012vy ; Cisterna:2014nua ; Feng:2015oea ; Sotiriou:2013qea ; Miao:2016aol ; Kuang:2016edj ; Babichev:2016rlq ; Benkel:2016rlz ; Filios:2018xvy ; Cisterna:2018hzf ; Giusti:2021sku ; Cisterna:2015uya ; Anabalon:2013oea ; Fontana:2018fof and the time-dependent scalar field Babichev:2013cya ; Babichev:2017lmw ; BenAchour:2018dap ; Takahashi:2019oxz ; Minamitsuji:2019shy ; Arkani-Hamed:2003juy . The hairy solution with a scalar hair in linear time dependence was initially ruled out due to instability Khoury:2020aya . However, another novel solution with a time-dependent scalar hair was recently constructed Bakopoulos:2023fmv and found to be stable Sirera:2024ghv . The stability conditions for a similar time-dependent background solution beyond Horndeski theories was also addressed in Mironov:2024yqa . Moreover, it was demonstrated in Hui:2012qt , that the no-hair theorem cannot hold when a Galileon field is coupled to gravity, but the static spherical solution only admits a trivial Galileon profile.
@@
An interesting new development in the examination of no-hair theorem in Horndeski theories is that the shift-symmetric Horndeski theory and beyond with the action Babichev:2017guv
@@
allow static and asymptotically flat black holes with a nontrivial static scalar field. In the above action, $g$ is the determinant of the metric and $R$ is the Ricci scalar. $\chi=-\partial^{\mu}\phi\partial_{\mu}\phi/2$ is the canonical kinetic term, $Q_{i}~{}(i=2,3,4,5)$ are arbitrary functions of the scalar field $\phi$ and the kinetic term $\chi$, and $Q_{i,\chi}\equiv\partial Q_{i}/\partial\chi$. In particular, a static hairy black hole in a specific quartic Horndeski theory, which includes an additional $\log$-term deforming from the standard Schwarzschild black hole, has been constructed in Bergliaffa:2021diw . This Horndeski hairy black hole soon attracts plenty of attention, and various theoretical and observational phenomena have been disclosed, for instances, thermodynamic and weak gravitational lensing Walia:2021emv ; Atamurotov:2022slw , the strong gravitational lensing Kumar:2021cyl , shadow constraint from EHT observation Afrin:2021wlj ; Vagnozzi:2022moj , superradiant properties in its rotating counterpart Jha:2022tdl ; Lei:2023wlt , photon rings in the black hole image Wang:2023vcv ; Hu:2023pyd and precessing and periodic orbits Lin:2023rmo . These studies disclosed the novel properties brought by the Horndeski hair, differentiated from the Schwarzschild black hole in GR, which are helpful for understanding the gravitational structure of the specific theory.
@@
Another significant question concerning this Horndeski hairy black hole is its stability. It was addressed in Walia:2021emv that the Horndeski hairy black hole is thermodynamically preferred. Later, we also demonstrated its dynamical stability under various external massless perturbed fields including the scalar field, electromagnetic field and Dirac field in Yang:2023lcm . However, up to now, its dynamical stability has not been completely investigated. To promote this topic, we have to perturb the metric itself and study the linear mode as a first important step, in which the modes of a spherically symmetric black holes usually can be classified into the even-parity (polar) modes and odd-parity (axial) modes.
@@
On the other hand, considering that the gravitational waves radiated during black hole coalescence encode rich information about the nature of spacetime, it is significant to study the gravitational waves signatures of such hairy black hole. In particular, the ringdown wave, as the associated gravitational emission at the late time during binary black hole coalescence, is expected to be modeled by the QNM frequencies. At this stage, the QNMs could be the key observable in constructing the patterns of such signal and therefore could be used to examine the no-hair theorem due to its highly dependence on the black hole parameter. Thus, the first step in this direction is to study their QNMs, which describe the ringdown stage after the binary black holes merger Berti:2015itd . The QNMs are the eigenmodes of a dissipative system, which inherently reflect the effects of all the system’s parameters. Consequently, if a black hole acquires an additional Horndeski hair, its QNMs will naturally encode the characteristics of this scalar hair. Due to the dissipative nature of the system, the frequencies of QNMs are typically composed of two parts: the real part and the imaginary part, the former gives the frequency of vibration while the latter describes the damping timescale. Readers can refer to the reviews Nollert:1999ji ; Berti:2009kk ; Konoplya:2011qq for more details.
@@
Thus, motivated by the above aspects, in this paper we will consider the dynamics of the gravitational odd-parity perturbations of a specific Horndeski hairy black hole. We expect that it has a simpler spectrum than the even-parity modes because it decouples from the perturbations of the scalar field, in contrast to the even-parity modes. We shall study the time-domain evolution and QNM frequencies for the odd-parity perturbations, then further answer the question whether the hairy black hole is stable under such perturbations. Moreover, in order to check the imprint of Horndeski hair on the GW in ringdown stage and give some insight into the test of no-hair theorem, we will use the yielding QNM data to construct the ringdown GW waveform and further calculate the Fisher information matrix (FIM) Berti:2005ys to obtain the measurement error of the detection of the Horndeski hair.
@@
This paper is organized as follows. In section II, we derive the modified Regge-Wheeler equation of of the gravitational odd perturbation, and give a proposal to refine it in a specific Horndeski hairy black hole. In section III, we numerically solve the master equations, and extract the QNM frequencies of the gravitational odd perturbation. We analyze the effects of scalar hair $Q$, the angular number $\ell$, and the overtones $n$ on the QNM frequencies. In section IV, we first briefly review the waveform modeling in the ringdown stage, and then investigate the detectability of the specific Horndeski hair by calculating the error bound in cases with single-mode and mixed-mode estimations. Section V concludes with our key findings and outlooks for future work. Throughout the paper, we will set $c=G=1$ unless we restate it.
@@ II.1 Review on the action and Horndeski hairy black hole
Horndeski gravity is the most general scalar-tensor field theory in four dimensions which possesses second-order field equations Horndeski:1974wa . In this subsection, we shall focus on a special subclass of Horndeski theory and review the static hairy black hole constructed in Bergliaffa:2021diw . The action is
@@
and then solving the field equation gives a non-stealth black hole solution
@@
where $M$ is the integral constant related to the black hole mass and $Q$ is fixed by the parameters of the theory. This indicates that the scalar hair should be classified as a secondary hair rather than a primary scalar hair, the latter requires $Q$ to be an integration constant appearing in the metric Bakopoulos:2023fmv . It is noted that $\alpha_{42}$ and $\alpha_{22}$ should have opposite signs in (10) to admit a real scalar field.
Thus, the subclass Horndeski theory with the action (2) with some specific settings admits a Horndeski hairy black hole with the metric Bergliaffa:2021diw
@@
in which the hairy term appears as a $\log$ term deforming the standard Schwarzschild black hole. $M$ and $Q$ are related to the mass and hairy charge of the black hole. It is obvious that the metric is asymptotically flat since $f(r)\mid_{r\to\infty}=1$. Let us briefly analyze the structure of the hairy metric (11). For arbitrary $Q$, $r=0$ represents an intrinsic singularity as the curvature scalar is singular, and $f(r)=0$ always admits the solution $r_{+}=2M$. In addition, (i) when $Q\to 0$, the metric reduces to the Schwarzschild case. (ii) When $Q>0$, $r_{+}=2M$ is the unique root of $f(r)=0$, so it has a unique horizon, i.e., the event horizon for the hairy black hole which resembles Schwarzschild black hole. (iii) For $-2M<Q<0$, $f(r)=0$ has two distinguishable positive real roots: the larger one is $r_{+}=2M$ denoting the position of event horizon, and the smaller one $r_{-}$ is the Cauchy horizon, lying inside the event horizon. The Cauchy horizon $r_{-}$ increases as $Q$ decreases, and finally coincides with $r_{+}$ when $Q=-2M$, indicating the extreme case.
@@ II.2 Odd-parity perturbation and modified Regge-Wheeler equation
In this subsection, we will consider the gravitational perturbation to the Horndeski hairy black hole (11). We will focus on the odd-parity perturbation at linear order, thus, we can ignore the coupling between the even-parity perturbation and the perturbation of the energy-momentum tensor from the scalar field, both of which decouple from the odd-parity perturbation. The gravitational perturbation of general spherically symmetric black holes in Horndeski gravity was well addressed in Kobayashi:2012kh ; Kobayashi:2014wsa , based on which we shall review the derivation of the modified Regge-Wheeler equation and present our refined version of this equation.
Under the Regge-Wheeler gauge Regge:1957td , the metric perturbed by the odd-parity mode takes the form
$g_{\mu\nu}$ $=g^{background}_{\mu\nu}+h_{\mu\nu,\ell m}^{\text{odd}}\qquad\qquad\qquad\quad$ (12)
$=\begin{pmatrix}A(r)&0&0&0\\ 0&B(r)&0&0\\ 0&0&C(r)&0\\ 0&0&0&C(r)\sin^{2}{\theta}\end{pmatrix}$ $+\begin{pmatrix}0&0&0&h_{0}(r,t)\\ 0&0&0&h_{1}(r,t)\\ 0&0&0&0\\ h_{0}(r,t)&h_{1}(r,t)&0&0\end{pmatrix}\sin{\theta}\partial_{\theta}Y_{\ell m}$
where the spherical harmonics can be replaced by Legendre polynomials by setting the azmuthal number $m=0$ without loss of generality, i.e. $Y_{\ell m}\mid_{m=0}=\sqrt{\frac{2\ell+1}{4\pi}}P_{\ell}(\cos\theta)$, because the background hairy metric is spherically symmetric. By substituting the perturbed metric (12) into the action (2), we can obtain a second order on-shell action as